### v1

# RGB-to-Polarization Estimation: A New Task and Benchmark Study

Abstract
Polarization images provide rich physical information that is fundamentally absent from standard RGB images, benefiting a wide range of computer vision applications such as reflection separation and material classification. However, the acquisition of polarization images typically requires additional optical components, which increases both the cost and the complexity of the applications. To bridge this gap, we introduce a new task: RGB-to-polarization image estimation, which aims to infer polarization information directly from RGB images. In this work, we establish the first comprehensive benchmark for this task by leveraging existing polarization datasets and evaluating a diverse set of state-of-the-art deep learning models, including both restoration-oriented and generative architectures. Through extensive quantitative and qualitative analysis, our benchmark not only establishes the current performance ceiling of RGB-to-polarization estimation, but also systematically reveals the respective strengths and limitations of different model families — such as direct reconstruction versus generative synthesis, and task-specific training versus large-scale pre-training. In addition, we provide some potential directions for future research on polarization estimation. This benchmark is intended to serve as a foundational resource to facilitate the design and evaluation of future methods for polarization estimation from standard RGB inputs.
1 Introduction
Polarization images contain rich physical information that is not captured by standard RGB cameras, such as birefringence, surface stress, roughness, and other material properties wolff1990polarization . This polarization information provides valuable visual cues that enhance a variety of computer vision tasks, such as reflection separation nayar1997separation ; lei2020polarized ; pang2022progressive , material analysis tasks like classification and segmentation wolff1990polarization ; mei2022glass ; liang2022multimodal , and shadow removal zhou2025polarization .
Figure 1: Comparison between sensor-based methods and our polarization estimation approach. Conventional methods rely on physical acquisition systems (e.g., polarization cameras or rotating polarizers), whereas our method leverages RGB inputs and neural networks to estimate polarization information without requiring dedicated hardware. The predicted polarization images can be readily applied to a variety of downstream tasks.
However, polarization images are not widely accessible in practice, since their acquisition requires specialized hardware, such as polarization cameras or standard cameras equipped with a rotating polarizer (see Figure 1). These devices are often expensive and inconvenient to operate, making polarization imaging impractical for widespread or everyday use. While polarization data remains difficult to obtain, standard RGB images are widely available. This raises a fundamental question: Can polarization information be estimated directly from RGB inputs without relying on dedicated polarization sensors?
In response to this question, we introduce a new task: RGB-to-polarization image estimation. As illustrated in Figure 1, the goal is to leverage neural networks to estimate polarization information directly from RGB inputs. To the best of our knowledge, this is the first work that explores sensor-free polarimetric imaging using neural networks. In this work, polarization information is represented using the Stokes parameters. We treat the RGB image—corresponding to the total intensity ($\mathrm{S}_{0}$)—as input, and train models to estimate the remaining polarization components, $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$, which respectively characterize horizontal/vertical linear polarization ($0^{\circ}$/$90^{\circ}$), diagonal linear polarization (±$\pm 45^{\circ}$), and circular polarization. These components are essential for a wide range of downstream polarization-based vision tasks.
Building on this task, we establish a comprehensive benchmark for RGB-to-polarization image estimation. Leveraging one of the latest RGB-polarization datasets jeon2024spectral , we standardize evaluation protocols to ensure consistent and comparable results across methods. Multiple deep learning models, spanning both restoration-based and generative architectures, are further evaluated to assess their effectiveness on this task. Through systematic analysis, our benchmark reveals the strengths and limitations of existing approaches and offers insights for future development in sensor-free polarization estimation. Our contributions are summarized as follows:
  • Task and Benchmark: We introduce RGB-to-polarization estimation as a new computer vision task and establish the first benchmark for sensor-free polarization imaging. Built upon a recent RGB-polarization dataset, the benchmark features standardized evaluation protocols that enable consistent and reproducible comparisons across methods.
  • Comparative Study: We evaluate a diverse set of deep learning models, including both restoration-based and generative approaches, to assess their performance on this task and uncover key limitations.
  • Insights for Future Research: Our analysis provides practical guidance for future model design and highlights open challenges in estimating polarization information from RGB images.
2 Related work
Polarization data
In the field of optics, light is typically characterized by four fundamental properties: amplitude, wavelength, phase, and polarization boyd2008nonlinear . Among these, polarization plays a pivotal role in light–matter interactions, providing discriminative cues that help distinguish surface characteristics such as smoothness, refractive index, and coating wolff1990polarization .
Several mathematical models have been developed to describe the polarization state of light. The Jones vector sheppard2014jones ; gori2001polarization describes light using two complex amplitudes for orthogonal electric field components. However, it is limited to fully polarized light and cannot account for partial or incoherent polarization, which are common in natural scenes. The Mueller matrix leroybrehonnet1997mueller ; gil2022polarized models the effect of materials on polarization, including depolarization, using a $4{\times}4$ real-valued matrix. While it accommodates both fully and partially polarized light, it demands extensive measurements and introduces 16 parameters, making it both experimentally demanding and computationally heavy. The polarimetric BRDFpriest2000polarimetric models material-specific polarized reflectance but requires dense angular sampling. In this work, we adopt the Stokes vector representation for its ability to comprehensively describe various polarization states and its applicability to incoherent light analysis schaefer2007measuring ; rubin2019matrix ; mcmaster1954polarization ; deboer2002review . A detailed explanation of how polarization is encoded in the Stokes parameters can be found in section 3.
Polarization datasets
Several polarization datasets are developed to support the analysis of polarization and spectral properties, as well as downstream vision tasks. Early works ba2020deep ; dave2022pandora ; liang2022material ; qiu2021demos collect small-scale data using trichromatic linear-polarization cameras, typically focusing on a limited number of objects or controlled indoor scenes. More recent efforts expand to larger-scale datasets designed for specific computer vision applications, such as reflection separation lei2020refl ; lyu2019refl and transparent object segmentation mei2022glass . Fan et al. fan2023fullst present the first full-Stokes dataset that includes both linear and circular polarization components. However, the dataset consists of only 64 flat objects, which limits its applicability to real-world scenarios. Jeon et al. jeon2024spectral propose a large-scale dataset that includes both trichromatic and hyperspectral Stokes measurements, covering more than 2,000 natural scenes under diverse illumination conditions. This dataset serves as the foundation for our benchmark, enabling RGB-to-polarization estimation in realistic and varied settings.
Restoration backbones
Existing image restoration backbones liang2021swinir ; wang2022uformer ; zamir2022restormer ; chen2022simple ; li2023efficient ; zamir2021multi are widely used in low-level vision tasks and can be naturally adapted for RGB-to-polarization image estimation. Transformer-based architectures such as SwinIR liang2021swinir , Uformer wang2022uformer , and Restormer zamir2022restormer leverage self-attention mechanisms to capture long-range dependencies, achieving strong performance across denoising, deraining, and deblurring tasks. In our benchmark, we adopt Uformer and Restormer as representative restoration backbones due to their effectiveness and generalizability.
In parallel, large-scale vision transformers pre-trained in a self-supervised manner, such as DINO caron2021emerging and MAE he2022masked , have shown strong representation learning capabilities across a range of tasks. These models are trained on massive image collections without manual labels, and their pre-trained weights can be transferred to downstream problems with limited supervision. In our setting, we explore whether the rich semantic priors learned by MAE can facilitate RGB-to-polarization estimation. Specifically, we initialize the MAE encoder with pre-trained weights and fine-tune the entire network for RGB-to-polarization estimation.
Generation backbones
Existing generative models, such as GANs goodfellow2014generative and diffusion models ho2020ddpm , have demonstrated strong capabilities in modeling complex image distributions. In this work, we explore whether diffusion models can be used to estimate polarization images from RGB inputs. Unlike restoration-based models such as Restormer, which perform direct pixel-wise regression, diffusion models learn to approximate the underlying data distribution and generate outputs through iterative denoising. This generative formulation allows them to capture uncertainty and produce more flexible predictions. Following our task setting, we adopt two representative diffusion models, WDiff ozdenizci2023restoring and DiT peebles2023scalable , both trained in a conditional setting where the RGB image guides the generation of polarization components.
In addition to training task-specific diffusion models from scratch, Stable Diffusion rombach2022high , a large-scale pre-trained model, is leveraged to estimate polarization images. Previous works wang2024exploiting ; lin2025seeing ; wu2024one show that the priors of pre-trained diffusion models can be effectively transferred to downstream tasks. We adopt two representative models, RealFill tang2024realfill and Img2ImgTurbo parmar2024one , which are originally designed for image inpainting and translation. These models are further fine-tuned for the RGB-to-polarization estimation task, enabling the synthesis of polarization components guided by RGB context.
Polarization-related tasks
When extended to computer vision tasks, polarization plays a critical role in distinguishing surface properties. Specifically, the polarization state of reflected light changes depending on factors such as a material’s refractive index, surface roughness, coating, or texture—enabling the classification of materials and object recognition wolff1990polarization . In addition, specular reflections can be suppressed by exploiting polarization properties, and the contrast of transparent or translucent objects can be enhanced using a polarizer riviere2017polarization . Furthermore, materials exhibiting birefringence or optical activity cause characteristic changes in polarization, often reflected in the $\mathrm{S}_{2}$ and $\mathrm{S}_{3}$ components, which are useful for biomedical imaging, tissue analysis, and fiber inspection deboer2002review . Finally, surface normals affect the polarization state of reflected light, which in turn provides geometric cues for shape or depth inference in tasks such as Shape-from-Polarization (SfP) based 3D reconstruction kadambi2015polarized . These cues are absent in standard RGB images, highlighting the unique and valuable role of polarization information.
(a) Poincaré sphere
(b) $\mathbf{S}=(\pm 1,0,0)$
(c) $\mathbf{S}=(0,\pm 1,0)$
(d) $\mathbf{S}=(0,0,\pm 1)$
Figure 2: (a) The Poincaré sphere serves as a geometric representation for describing all possible states of polarization using the Stokes parameters. (b–d) Polarization describes how the electric field of a light wave oscillates within the plane perpendicular to the direction of propagation. (b) Linear polarizations at 0° and 90°. (c) Linear polarizations at ±$\pm 45^{\circ}$ angles. (d) Right- and left-handed circular polarizations.
3 Method
In this section, we first provide a brief overview of the Stokes parameters that characterize the polarization state of the light, followed by the formulation of the RGB-to-polarization estimation task and the construction of our benchmark, including dataset preparation, evaluation protocols, and model evaluation.
3.1 Stokes parameters
The polarization state of light characterizes how the electric field oscillates within the plane perpendicular to the direction of propagation. In this work, we represent the polarization state of light using the Stokes vector and visualize it on the Poincaré sphere, as illustrated in Figure 2, where each point corresponds to a unique polarization state. A light wave’s polarization can be described using the four Stokes parameters $(\mathrm{S}_{0},\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})$, where $\mathrm{S}_{0}$ denotes the total intensity of the beam, and the remaining components collectively characterize the polarization state. Specifically, $\mathrm{S}_{1}$ represents the difference in intensity between 0 and 90 linear polarization; $\mathrm{S}_{2}$ corresponds to the difference between +45 and -45 linear polarization; and $\mathrm{S}_{3}$ indicates the difference between right- and left-handed circular polarization mcmaster1954polarization ; schaefer2007measuring .
Mathematically, the Stokes parameters can be expressed in terms of the polarization azimuth angle $\psi$ and the ellipticity angle $\chi$ as:
$\mathrm{S}_{1}$ $=\mathrm{S}_{0}\cos(2\psi)\cos(2\chi)$ (1)
$\mathrm{S}_{2}$ $=\mathrm{S}_{0}\sin(2\psi)\cos(2\chi)$ (2)
$\mathrm{S}_{3}$ $=\mathrm{S}_{0}\sin(2\chi)$ (3)
where $\mathrm{S}_{0}$ denotes the total light intensity, represented as an RGB image. On the Poincaré sphere, the polarization direction is represented by the vector $\mathbf{S}=(\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3})/\mathrm{S}_{0}$. For instance, $\mathbf{S}=(1,0,0)$ corresponds to 0 linear polarization, $(-1,0,0)$ to 90 linear polarization, $(0,1,0)$ and $(0,-1,0)$ to ±45 linear polarization, and $(0,0,\pm 1)$ to right- or left-handed circular polarization. Elliptical polarization states lie between these extremes, depending on the value of $\mathrm{S}_{3}$. Each component of the Stokes vector carries distinct material-sensitive information: high $\mathrm{S}_{1}$ values often indicate smooth dielectric or metallic surfaces due to dominant specular reflection; elevated $\mathrm{S}_{2}$ suggests birefringent or fibrous materials; and strong $\mathrm{S}_{3}$ values arise in scattering or optically active media, such as biological tissues or rough dielectrics riviere2017polarization .
3.2 Benchmark Design
Task definition
We define RGB-to-polarization image estimation as a pixel-wise prediction task, where the goal is to estimate polarization information from a single RGB input image. Given an RGB image $\mathbf{I}_{\text{RGB}}\in\mathbb{R}^{H\times W\times 3}$, where $H\times W$ denotes the spatial resolution, the objective is to estimate Stokes components $\mathbf{S}\in\mathbb{R}^{H\times W\times 9}$. Each Stokes component, $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$, is represented as a 3-channel image. The final output is obtained by concatenating these components along the channel dimension, resulting in $\mathbf{S}=[\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3}]$.
Dataset
Next, we build our benchmark on the recent large-scale RGB-polarization dataset proposed by Jeon et al. jeon2024spectral . It provides high-quality, spatially aligned image pairs consisting of RGB inputs $\mathrm{S}_{0}$ and their corresponding Stokes components $[\mathrm{S}_{1},\mathrm{S}_{2},\mathrm{S}_{3}]$, enabling supervised training. All images are resized to a fixed resolution of $H\times W$ before training. Both the RGB inputs and the Stokes targets are normalized to the range $[0,1]$ for consistent training across models.
Evaluation metrics
Following standard practices in low-level vision, we evaluate model performance using three complementary metrics: peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS). PSNR and SSIM assess pixel-level accuracy and structural fidelity, while LPIPS captures perceptual similarity in the feature space. All metrics are computed independently for each Stokes component and then averaged to report the overall performance.
3.3 Baselines
With the defined benchmark, we evaluate a range of representative deep learning models that fall into two categories: restoration-based and generation-based approaches.
Restoration-based approaches.
Two representative restoration backbones, Restormer zamir2022restormer and Uformer wang2022uformer , are selected for evaluation. Both models are originally designed for image-to-image restoration tasks, where the input and output are 3-channel RGB images. They have demonstrated strong performance on tasks such as denoising, deraining, and deblurring, making them suitable baselines for pixel-wise polarization prediction. In our implementation, we modify the output layer to produce 9 channels, corresponding to the concatenated Stokes components $\mathbf{S}\in\mathbb{R}^{H\times W\times 9}$ for supervision. The models are trained using an $L_{1}$ loss between the predicted and ground-truth Stokes components.
Then we further evaluate Masked Autoencoder (MAE) he2022masked , a vision transformer pre-trained using self-supervised learning on large-scale image datasets. MAE learns strong visual representations by reconstructing randomly masked image patches. With the pre-trained parameters, it can extract robust features for a variety of downstream tasks. In our setting, we modify the output channels of the final predictor projection layer to 9, corresponding to the concatenated Stokes components $\mathbf{S}\in\mathbb{R}^{H\times W\times 9}$ used for supervision. Both the MAE encoder and decoder are initialized with pre-trained weights. The predictor projection layer is initialized by inflating the original parameters through channel-wise replication to match the 9-channel output. The entire model is then fine-tuned end-to-end using an $L_{1}$ loss between the predicted and ground-truth Stokes components.
Generation-based approaches.
Next, we explore generative diffusion models for RGB-to-polarization estimation. Two types of models are evaluated: (1) task-specific diffusion models trained from scratch, and (2) large-scale pre-trained models adapted to our task.
For the first type, we adopt WDiff ozdenizci2023restoring and DiT peebles2023scalable , two representative conditional diffusion models. Both are trained to iteratively denoise a latent polarization representation conditioned on the input RGB image. In our setting, we treat the RGB image as the conditioning input and train the diffusion process to generate the 9-channel Stokes output $\mathbf{S}\in\mathbb{R}^{H\times W\times 9}$. The training objective minimizes the denoising error over the diffusion steps using standard diffusion loss formulations.
For the second type, we adapt pre-trained diffusion models to the polarization estimation task. Specifically, we fine-tune RealFill tang2024realfill and Img2ImgTurbo parmar2024one , which are originally developed for image inpainting and translation. These models are built upon Stable Diffusion rombach2022high and leverage rich visual priors learned from large-scale image-text datasets. Instead of modifying the model architecture, we train a separate model for each Stokes component, i.e., $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$, where each component is represented as a 3-channel image. This allows us to directly reuse the pre-trained UNet decoder without altering the output layer. Each model is fine-tuned in a conditional generation setting, where the RGB image serves as guidance. During training, we follow the default configurations and hyperparameters provided in the original implementations.
Table 1: Quantitative results on RGB-to-polarization image estimation. We report PSNR, SSIM, and LPIPS for each estimated Stokes component ($\mathrm{S}_{1}$, $\mathrm{S}_{2}$, $\mathrm{S}_{3}$), as well as their averages. Higher PSNR and SSIM and lower LPIPS indicate better performance. We highlight the best and second-best results for each component and the averages.
Method $\bm{\mathrm{S}_{1}}$ $\bm{\mathrm{S}_{2}}$ $\bm{\mathrm{S}_{3}}$ Average PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS Wdiff ozdenizci2023restoring 10.62 0.6454 0.4617 14.65 0.7172 0.2817 14.05 0.6840 0.3882 13.11 0.6822 0.3772 DiT peebles2023scalable 20.96 0.7984 0.3238 24.01 0.8607 0.1680 25.02 0.8636 0.2424 23.33 0.8409 0.2447 Realfill tang2024realfill 20.43 0.8194 0.2964 21.92 0.7573 0.2292 23.07 0.8308 0.2707 21.81 0.8025 0.2654 Img2ImgTurbo parmar2024one 21.47 0.8532 0.3931 23.65 0.8522 0.3465 24.87 0.9122 0.3230 23.33 0.8725 0.3542 Restormer zamir2022restormer 22.54 0.8495 0.3072 24.42 0.8764 0.1693 24.99 0.8932 0.2341 23.98 0.8730 0.2369 Uformer wang2022uformer 22.61 0.8482 0.3007 24.72 0.8721 0.1624 25.69 0.8963 0.2170 24.34 0.8722 0.2267 MAE he2022masked 22.73 0.8690 0.3521 25.54 0.8759 0.2194 25.94 0.9179 0.2338 24.74 0.8876 0.2684
4 Experiments
In our experiments, we use the dataset provided by Jeon et al. jeon2024spectral for training and evaluation. Specifically, the first 1,000 RGB–Stokes image pairs are used for training, and the last 200 pairs from the dataset are reserved for testing. For a fair comparison, all image pairs are resized to 256$\times$256 for training and testing across all baseline models.
4.1 Implementation details
All experiments are conducted on a server equipped with four NVIDIA RTX A5000 GPUs, each with 24 GB of memory. To ensure reproducibility, we will release all code, model checkpoints, and configuration files upon acceptance. The following parts detail the network architectures and training settings used across all evaluated models.
Restoration baselines
Restormer zamir2022restormer consists of four hierarchical levels, each containing a series of transformer blocks. The numbers of transformer blocks in the four levels are set to 4, 6, 6, and 8, respectively. The embedding dimension is set to 48. Uformer wang2022uformer follows an encoder–decoder architecture with four encoder layers and four decoder layers. Each layer contains two transformer blocks, and the embedding dimension is set to 32. MAE he2022masked adopts an asymmetric encoder–decoder design, where the encoder and decoder contain 24 and 8 transformer blocks, respectively.
The training steps for the three networks are set to 70,000, 70,000, and 40,000, respectively. Due to differences in patch size and network depth, we adopt different batch sizes for training. For each model, we use the maximum batch size that fits within GPU memory constraints. The batch sizes are set to 16, 48, and 128, respectively. During training, we use the Adam optimizer with an initial learning rate of 1.5$\times$10-4.
Figure 3: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 4: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Generation baselines
WDiff ozdenizci2023restoring adopts a hierarchical U-Net architecture with six resolution levels, each containing two residual blocks. Attention is applied at the $16\times 16$ resolution. DiT peebles2023scalable is a Vision Transformer-based diffusion model with 10 transformer blocks, each using 6 attention heads and a hidden size of 768. RealFill tang2024realfill is built on top of Stable Diffusion, using a U-Net backbone with cross-attention. It incorporates Low-Rank Adaptation (LoRA) modules with rank 8 and dropout 0.1 to fine-tune both the UNet and the text encoder. Only LoRA parameters are updated during training, while the base model remains frozen. Img2ImgTurbo parmar2024one also builds on Stable Diffusion, but improves generation efficiency by directly injecting the RGB latent into the denoising UNet. It integrates LoRA adapters into both the UNet and VAE components, with LoRA ranks set to 8 and 4, respectively, enabling lightweight and flexible fine-tuning.
The training steps for WDiff, DiT, RealFill, and Img2ImgTurbo are set to 90,000, 90,000, 10,000, and 8,000, respectively. We use the Adam optimizer with an initial learning rate of 1.5$\times$10-4 for WDiff and DiT, and adopt batch sizes of 128 and 258, respectively. For RealFill and Img2ImgTurbo, due to higher memory consumption, the batch sizes are limited to 12 and 4, respectively. To simulate a larger effective batch size, we apply gradient accumulation with a factor of 4 for Img2ImgTurbo. Following tang2024realfill , we adopt separate learning rates for LoRA modules in RealFill: 2e-4 for the UNet and 4e-5 for the text encoder. For Img2ImgTurbo parmar2024one , we set a unified learning rate of 5e-6 for all learnable parameters.
4.2 Quantitative evaluation
Table 1 shows the quantitative results for RGB-to-polarization estimation across all baseline models. For each method, we evaluate the predicted Stokes components ($\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$) using three commonly used metrics: PSNR, SSIM, and LPIPS. Among all methods, MAE he2022masked achieves the best overall performance, attaining the highest average PSNR (24.74), SSIM (0.8876), and a competitive LPIPS (0.2684). While Restormer zamir2022restormer and Uformer wang2022uformer are trained from scratch, both models also exhibit strong performance, particularly in structural fidelity and perceptual similarity. Notably, Uformer achieves the lowest average LPIPS score of 0.2267 across all evaluated methods.
Diffusion-based models exhibit inconsistent performance across different architectures. WDiff ozdenizci2023restoring shows relatively low reconstruction quality, while DiT peebles2023scalable demonstrates notable improvements across all metrics. Among the pre-trained diffusion models, Img2ImgTurbo parmar2024one consistently outperforms RealFill tang2024realfill on most metrics, particularly in PSNR and SSIM. These results establish a clear performance landscape across different model families and provide a foundation for further comparative analysis.
Figure 5: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 6: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
4.3 Qualitative evaluation
Figures 3, 4, 5, and 6 show qualitative results of the estimated Stokes components generated by DiT peebles2023scalable , Img2ImgTurbo parmar2024one , Uformer wang2022uformer , and MAE he2022masked . Each column visualizes the reconstructed $\mathrm{S}_{1}$, $\mathrm{S}_{2}$, and $\mathrm{S}_{3}$ components produced by each method. The results reveal diverse reconstruction characteristics across models, including differences in texture sharpness, structural consistency, and polarization pattern styles. Overall, MAE and Uformer demonstrate superior visual quality compared to diffusion-based approaches. MAE is particularly effective in preserving fine structural details and material boundaries, with minimal artifacts across all Stokes components. However, the restored details from restoration-based methods still deviate from the correct results, indicating that such models may still lack the capacity to capture the underlying polarization cues precisely. In contrast, diffusion-based methods struggle to reconstruct the correct appearance and structure of polarization components, indicating difficulty in modeling these signals. Although Img2ImgTurbo achieves higher PSNR scores than DiT, its visual results are less faithful to the ground truth, particularly in terms of structural detail. This suggests a trade-off between pixel-level accuracy and perceptual quality among diffusion-based models.
These qualitative observations highlight that both restoration-based and generative-based methods have limitations in RGB-to-polarization estimation. Restoration models may lack fine polarization accuracy, while generative models often struggle with structural consistency.
4.4 Discussion
Restoration vs. Generation
Table 1 reveals a clear performance gap between restoration-based and generation-based approaches. Restoration models, such as Restormer and Uformer, consistently achieve higher PSNR and SSIM scores than diffusion-based models like WDiff and DiT. Among the pre-trained baselines, MAE also outperforms both Img2ImgTurbo and RealFill by a substantial margin across most metrics. These results suggest that the restoration backbones are more effective for Stokes component prediction, likely due to their capacity for precise pixel-level estimation.
Although WDiff struggles to produce accurate reconstructions, more advanced diffusion models such as DiT demonstrate marked improvements across all metrics. This indicates that task-specific diffusion models still hold promise for polarization estimation, particularly with further architectural or training enhancements.
Pre-training vs. Training from scratch
Pre-trained models exhibit clear advantages in RGB-to-polarization estimation. MAE achieves the highest average PSNR and SSIM, outperforming Uformer by a noticeable margin, demonstrating the effectiveness of transferring rich visual representations learned from large-scale RGB data. Img2ImgTurbo also outperforms task-specific diffusion models like WDiff and DiT by leveraging the strong generative prior of Stable Diffusion. In contrast, RealFill fine-tunes only the LoRA modules in the UNet while keeping the VAE frozen. This leads to information loss during latent encoding and decoding, which degrades the quality of polarization prediction.
One contributing factor to the improved performance is the intrinsic correlation between RGB images and polarization information. Since RGB images encode overall light intensity and structural patterns, models pre-trained on RGB data are better equipped to extract cues relevant for polarization estimation, even in the absence of direct supervision. These findings indicate that incorporating pre-trained weights, whether through self-supervised representation learning or large-scale generative modeling, is a promising strategy to enhance the accuracy and robustness of sensor-free polarization prediction.
Insights for future research
Our benchmark reveals several insights for future research. First, although both restoration-based and generation-based methods achieve promising results, there remains significant room for improvement in fine-grained detail reconstruction and polarization fidelity. Second, our benchmark leverages existing state-of-the-art backbones to establish strong baselines. However, future research may benefit from incorporating physical constraints, as the Stokes components inherently encode rich physical properties of light. Third, since acquiring large-scale paired data for RGB-to-polarization training is challenging, future work may explore self-supervised methods using unlabeled polarization data. Fourth, although our benchmark adapts image pre-trained models for polarization estimation, the adaptation is not specifically tailored to this task. Future research may explore more effective adaptation or fine-tuning methods to better capture polarization-related features. Finally, as the estimated polarization information may contain inaccuracies, it can affect downstream applications. Estimating polarization along with a confidence map is therefore an important direction for future research.
5 Conclusion
This paper introduces and benchmarks a new task: RGB-to-polarization image estimation, which aims to infer polarization information from standard RGB inputs without requiring specialized sensors. We formalize the task using Stokes parameters and construct the first comprehensive benchmark based on a recent large-scale dataset. Our evaluation covers various deep learning models, including restoration-based backbones and generation-based backbones. Through extensive quantitative and qualitative analysis, we reveal the strengths and limitations of existing approaches, providing a detailed performance landscape for this underexplored problem. Our results indicate that pre-trained models, such as MAE and Stable Diffusion with LoRA, offer strong prior knowledge that can be effectively transferred to polarization estimation, leading to consistently improved performance. We hope this benchmark will serve as a foundation for future research, fostering the development of accurate and efficient polarization estimation methods from RGB images alone.
Appendix A More results
We present supplementary qualitative results of the estimated Stokes components generated by DiT peebles2023scalable , Img2ImgTurbo parmar2024one , Uformer wang2022uformer , and MAE he2022masked . The results are illustrated in Figures 721. Among the methods, MAE and Uformer produce consistently better visual quality than the diffusion-based approaches. Nonetheless, there remains substantial room for improvement in accurately estimating polarization data.
Appendix B Limitations
Although our benchmark provides a comprehensive evaluation of RGB-to-polarization estimation, it still has several limitations due to practical constraints. First, our evaluation is based on the dataset provided by Jeon et al. jeon2024spectral . While it includes more than 2000 scenes, the diversity of materials and surface types is limited. As a result, polarization estimation under complex material properties, such as metals, birefringent materials, and biological tissues, remains underexplored. Second, adverse conditions such as weather-induced degradations or extreme noise can degrade image quality and thus affect polarization estimation. Our benchmark does not evaluate these cases due to the lack of such scenarios in existing datasets. Expanding the dataset to include more challenging and diverse environments is an important direction for future work.
Appendix C Societal impact
Polarization data has proven useful in many computer vision tasks, such as reflection separation and material classification. However, it remains difficult to obtain polarization data on consumer devices, limiting the practical use of polarization-based methods. This paper introduces a new task: RGB-to-polarization image estimation, which aims to infer polarization information directly from RGB images. By reducing reliance on specialized sensors, this approach has the potential to democratize access to polarization imaging and enable broader adoption in both scientific and industrial settings.
Figure 7: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 8: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 9: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 10: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 11: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 12: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 13: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 14: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 15: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 16: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 17: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 18: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 19: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 20: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .
Figure 21: Qualitative comparison of estimated polarization components from RGB input. Results are shown for Uformer wang2022uformer , MAE he2022masked , DiT peebles2023scalable , and Img2ImgTurbo parmar2024one .